OpenStudy (anonymous):

MEDAL AND FAN TO BEST ANSWER :D What does it mean to say that a number is a perfect square? Give one example of a number that is a perfect square and one that is not. Explain your examples. _define the term of a perfect square _provide examples, and explain why each is or is not a perfect square.

OpenStudy (johnweldon1993):

A perfect square is a number that you can take the square root of and get a whole number Example \[\large \sqrt{9} = 3\] so we know that 3 is a perfect square...however if we have \[\large \sqrt{5}\] is that a whole number? no..so 5 is NOT a perfect square in general, its easy to remember if you take EVERY number a square it...we will get a perfect square 1^2 = 1 2^2 = 4 3^2 = 9 4^2 = 16 1,4,9 and 16 are all examples of perfect squares

OpenStudy (dannyrod2000):

Examples of perfect squares 9 9 is a perfect square because it can be expressed as 3 * 3 (the product of two equal integers) 16 16 is a perfect square because it can be expressed as 4 * 4 (the product of two equal integers) 25 25 is a perfect square because it can be expressed as 5 * 5 (the product of two equal integers) Non examples of perfect squares 8 8 is a not perfect square because you cannot express it as the product of two equal integers 5 5 is a not perfect square because it cannot be expressed as the product of two equal integers 7 7 is a not perfect square because you cannot express it as the product of two equal integers